OpenStudy (babynini):

How do I integrate an equation where x is in the numerator and denominator? My equation is x/(sqrt(x^2+6)

OpenStudy (amistre64):

thats the derivative of the bottom?

OpenStudy (babynini):

@perl the last number is 6 :)

OpenStudy (amistre64):

there are a few things to look at when trying to determine the best process, the first thing i look at is if the top and bottom. since ln(u) derives to u'/u

OpenStudy (perl):

$$\large \int \Large \frac{x}{\sqrt{x^2+6}}$$

OpenStudy (babynini):

@amistre64 Could I do a U substitution? or, what should I do next?

OpenStudy (amistre64):

spose we wanted to take the deriavtive of: sqrt(x^2+a) what would it look like? and yes, a u sub would most likely show us this

OpenStudy (astrophysics):

Yes, do a u sub

OpenStudy (astrophysics):

\[\int\limits \frac{ x }{ \sqrt{x^2+6} }dx \implies u = x^2+6, du = 2x dx\] can you finish it off?

OpenStudy (amistre64):

or u = sqrt(x^2 + 6) :)

OpenStudy (amistre64):

du = 2x/2sqrt(x^2+6) dx

OpenStudy (astrophysics):

Yes hehe :)

OpenStudy (amistre64):

in some cases we have to decompose into seperate fractions, sometimes we can modify by adding 0 or multiplying by 1, other times its a trig function, other times .... but quite frankly these integrable functions are rare and they are trying to get you comfortable with the process, a false sense of security :)

OpenStudy (rational):

Haha like almost all polynomial equations have no algebraic solution because almost all polynomials have degree greater than 4 ;p

OpenStudy (babynini):

@Astrophysics Yeah, I can finish it off. Is the answer 1/3(x^2+6)(^3/2) ?

OpenStudy (astrophysics):

Mhm, not quite, \[\frac{ 1 }{ x^{n} } \implies x^{-n}\]

OpenStudy (astrophysics):

Oh wow it saved over my work.

OpenStudy (babynini):

oh i see! so u^1/2 or (2x+6)^1/2 is the correct answer?

OpenStudy (astrophysics):

\[\frac{ du }{ 2 }=xdx \implies \frac{ 1 }{ 2 } \int\limits \frac{ du}{ \sqrt{x} } \implies \frac{ 1 }{ 2 } \int\limits u^{-1/2} du\]

OpenStudy (astrophysics):

Yes, (2x+6)^(1/2) :)

OpenStudy (astrophysics):

+C

OpenStudy (babynini):

What! I get a C+? xD A+ for you, Astrophysics, you're the best. Thank you so much :)

rvc (rvc):

one error its root u :) @Astrophysics All the best @Babynini

OpenStudy (astrophysics):

Yes, nice catch, I was just redoing the work as I saved over my previous latex haha, thanks for noticing :)!

OpenStudy (astrophysics):

@Babynini Haha, +C after you're done integrating as it's a indefinite integral you must know there is a constant :)

OpenStudy (babynini):

oooh...thanks, forgot about that xP

OpenStudy (astrophysics):

No worries!